# Pressure-dependent electronic superlattice in the Kagome-superconductor CsV3Sb5

### v1

@@ Acknowledgements
We acknowledge fruitful discussions with R. Fernandes, R. Heid, M. Merz and S. M. Souliou. This research received support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), under projects 422213477 (TRR 288 project B03) and 247310070 (CRC 1143 project C06 and B05). T.R., F.S., C.F. and J.G. express gratitude to the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter (ct.qmat EXC-2147, Project No. 390858490) for financial backing. Additionally, we extend our appreciation to the European Synchrotron Radiation Facility (ESRF) for providing access to their synchrotron radiation facilities.

### v2

@@ Acknowledgements
We acknowledge fruitful discussions with R. Fernandes, R. Heid, M. Merz and S. M. Souliou. This research received support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), under projects 422213477 (TRR 288 project B03) and 247310070 (CRC 1143 project C06 and B05). T.R., F.S., C.F., and J.G. express gratitude to the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter (ct.qmat EXC-2147, Project No. 390858490) for financial backing. Additionally, we extend our appreciation to the European Synchrotron Radiation Facility (ESRF) for providing access to their synchrotron radiation facilities.